OpenStudy (anonymous):

what is the y intercept of y+3/5x+8/-5

OpenStudy (anonymous):

Is that y+(3/5)x EQUALS (8/-5)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = mx + c is the general form of the equation where m is the slope and c is the intercept now c = -8/5

OpenStudy (anonymous):

The y-intercept is the "b" would you have the equation in the form y=mx+b. For this example subtract the (3/5)x to the other side rendering. y=(-3/5)x+(8/-5) so (8/-5) is your "b" (i.e., your y-intercept)

OpenStudy (anonymous):

I TRIED THAT ALREADSY IT SAYS IT WRONG